OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. Choices... x intercept = (8, 0); y intercept = (0, 6) x intercept = (-8, 0); y intercept = (0, -6) x intercept = (-6, 0); y intercept = (0, 8) x intercept = (6, 0); y intercept = (0, -8) WHICH ONE?!

OpenStudy (anonymous):

If you convert that equation to to y = mx + b form(slope intercept form), it would be y = 4/3x - 8. Then when you graph it, the x and y intercept would be: x intercept = (6, 0); y intercept = (0, -8)

OpenStudy (anonymous):

4 option. if (a,0) and (0,b) are intercepts of a straight line then (X/a)+(Y/b)=1 will be the equation of the straight line. so by deducing you question to that form we get 4th option .